A population of values has a normal distribution with μ=65.1μ=65.1 and σ=4.8σ=4.8. You intend to draw a random sample of size n=49n=49.

Find P76, which is the mean separating the bottom 76% means from the top 24% means.
P76 (for sample means) =
Answer #1
mean μ= 65.1
standard deviation σ= 4.8000
sample size       =n= 49
std error=σ=σ/√n= 0.686
for 76th percentile critical value of z= 0.71
therefore corresponding value=mean+z*std deviation= 65.59